# Recurrent Deep Kernel Learning of Dynamical Systems

### v1

@@ 1 Introduction
A digital twin [1, 2, 3, 4] is a virtual replica that mimics the structure, context, and behaviour of a physical system. The digital twin is synchronized to its physical twin with data in real time to predict dynamical behaviours and inform critical decisions. Constructing a digital twin typically requires accurate modelling of complex physical phenomena that are often described by nonlinear, time-dependent, parametric partial differential equations (PDEs). Full-order models (FOMs) generated by detailed solvers are often computationally demanding and thus not suitable for the multi-query, real-time contexts in digital twinning, such as uncertainty quantification [5, 6, 7], optimal control [8, 9, 10], shape optimization [11], parameter estimation [12, 13], and model calibration [14, 15]. Therefore, constructing computationally efficient reduced-order models (ROMs) with controlled accuracy is crucial for dealing with real-world systems [16].
@@
Generally speaking, any surrogate model that reduces the computational cost of a FOM can be considered a ROM. ROMs aim to intelligently represent high-dimensional dynamical systems in carefully-established low-dimensional latent spaces, so that the low-dimensionality improves computational efficiency, and the accuracy remains under control [17, 18, 19, 20, 21, 22, 16]. Despite the wide variety of approaches that can be found in the literature, we can identify two major categories of ROM approaches: intrusive and non-intrusive methods. Intrusive ROM techniques require access to the full-order solvers, which is often inconvenient in industrial implementations, especially for the cases involving legacy code and/or readily executed software. To overcome this limitation, non-intrusive ROM techniques have been developed to learn low-dimensional representations primarily from data, without accessing the code of FOMs. Such flexibility of non-intrusive ROMs has recently motivated the development of a vast amount of data-driven methods, such as dynamic mode decomposition [23, 24], reduced-order operator inference [25, 26, 27, 28], sparse identification of reduced latent dynamics [29, 30, 31, 32, 33, 34], manifold learning using deep auto-encoders [35, 36, 37, 38], data-driven approximation of time-integration schemes [39], and Gaussian process modeling for low-dimensional representations [40, 41].
@@
However, learning ROMs from noisy data is extremely challenging. While many works in the literature have focused on learning compact representation of high-dimensional data, for example using neural networks (NNs) [42, 43], or quantifying uncertainties using, for example, Gaussian processes (GPs) [44], herein we argue that both challenges must be tackled simultaneously for discovering ROMs properly. NNs excel in learning complex and nonlinear dependencies of the data, while they tend to struggle when quantifying uncertainties. Conversely, GPs struggle with large dataset due to the limited expressivity of the kernels and the need for the inversion of the covariance matrix, while they excel to quantify uncertainties. To overcome these limitations, and to leverage on the positive features of both NNs and GPs synergistically, deep kernel learning (DKL) was introduced in recent years [45, 46]. DKL aims to combine the best of the GP and NN worlds by constructing a GP taking as input the data processed by a NN in order to learn expressive kernels that can model complex relations of the data.
@@
In this paper, we exploit the DKL idea to develop a non-intrusive ROM that can deal with both high-dimensional and noisy data. In particular, with reference to Figure 1, our method includes a DKL encoder to compress high-dimensional measurements into low-dimensional distributions of state variables, a recurrent DKL latent dynamical model to predict the system’s evolution over time, and a decoder to reconstruct the measurements and interpret the latent representations.
@@
Figure 1: Proposed framework for reduced-order modelling of dynamical systems. We encode the measurements $\mathbf{x}_{t}$ at different time instances into the latent variables $\mathbf{z}_{t}$ by means of a deep kernel learning encoder. Then, we feed a sequence of length $H$ of consecutive latent variables $\mathbf{z}_{t-H:t}$, actions $\mathbf{u}_{t-H:t}$, and parameters $\mathbf{p}$ to a recurrent deep kernel learning to predict the next latent variable $\mathbf{z}_{t+1}$. The measurements are then reconstructed $\hat{\mathbf{x}}$ by means of a decoder from the latent variables.
@@
The model is trained without labelled data in an unsupervised manner by only relying on high-dimensional and noisy measurements of the system. We show the capabilities of our framework in two challenging numerical examples, namely, the dynamics of (i) a double pendulum and of (ii) a distributed reaction-diffusion system, assuming in both cases to deal with measurements over a two-dimensional spatial region at different time instants – thus mimicking a set of observations acquired by a camera. Our contribution is threefold:
@@
  • we improve the prediction accuracy and consistency over long horizons of the DKL framework introduced in [41] by modelling the latent dynamics using a recurrent NN,
@@
The paper is organised as follows: in Section 2, we introduce the building blocks of our framework, namely GPs, and DKL, and in Section 3, we describe our novel DKL-based method. Section 4 presents the numerical experiments, shows and discusses the results, and Section 5 concludes the paper.
@@ 2.2 Deep Kernel Learning
The GP hyperparameters $\boldsymbol{\gamma}$ and the NN paramameters $\boldsymbol{\theta}$ are jointly learned by maximizing the log marginal likelihood (see Equation (3)). It is possible to use the chain rule to compute derivatives of the log marginal likelihood, that we indicate with $\mathcal{L}(\boldsymbol{\gamma},\boldsymbol{\theta})$, with respect to kernel hyperparameters $\boldsymbol{\gamma}$ and the NN parameters $\boldsymbol{\theta}$, thus obtaining:
@@
Optimizing the GP hyperparameters through Equation (3) requires to repeatedly invert the covariance matrix $k_{\boldsymbol{\gamma}}(X,X)+\sigma_{\epsilon}^{2}I$, with size $N\times N$. However, for large datasets, i.e., when $N\gg 1$, the full covariance matrix may be too large to be stored in memory and inverted. In these cases, it is common practice in machine learning to utilize minibatches of randomly-samples data points to compute the loss functions and compute the gradients. However, when we use minibatches to train the GPs, the posterior distribution is not Gaussian anymore, even when a Gaussian likelihood is used. When dealing with non-Gaussian posteriors, we can rely on variational inference (VI) [47]. VI is an approximation technique for dealing with non-Gaussian posteriors, deriving from minibatch training and/or non-Gaussian likelihoods. VI allows one to approximate the posterior by the best-fitting Gaussian distribution using a set of samples, often called inducing points, from the posterior.
@@ 3.2 Model Architecture
    a recurrent SVDKL forward dynamical model predicting the evolution of the system using a history of latent state variables $\mathbf{z}_{t-H},\cdots,\mathbf{z}_{t}$ [^2] For conciseness of the notation, we write $\mathbf{z}_{t-H},\cdots,\mathbf{z}_{t}=\mathbf{z}_{t-H:t}$ and $\mathbf{u}_{t-H},\cdots,\mathbf{u}_{t}=\mathbf{u}_{t-H:t}$. and control inputs $\mathbf{u}_{t-H},\cdots,\mathbf{u}_{t}$, and parameters $\mathbf{p}$, if available.
@@
With reference to Figure 3, the decoder $\psi$ is modeled using an NN with parameters $\boldsymbol{\theta}_{\psi}$. The decoder maps a sample $\mathbf{z}$ to the conditional distribution $p(\hat{\mathbf{x}}|\mathbf{z})$. Similar to VAEs [50], the distribution $p(\hat{\mathbf{x}}|\mathbf{z})$ is approximated to be independently distributed Gaussian distributions. In general, the decoding NN $\psi$ learns the means and variances of these distributions. Following [51, 52], however, our decoder in this work only learns the mean vector of $\hat{\mathbf{x}}|\mathbf{z}$, while the covariance is set to the identity matrix $I$, i.e., $\psi:\mathbb{R}^{|\mathbf{z}|}\rightarrow[0,1]^{|\mathbf{x}|}$ and
@@
With reference to Figure 4, the forward dynamical model $\xi:\mathbb{R}^{|\mathbf{z}|}\times\mathbb{R}^{|\mathbf{u}|}\times\mathbb{R}^{|\mathbf{p}|}\rightarrow[0,1]^{|\mathbf{z}|}$, with parameters $\boldsymbol{\theta}_{\xi}$ and hyperparameters $\boldsymbol{\gamma}_{\xi},\sigma_{\xi}^{2}$, is a SVDKL with a recurrent NN architecture, i.e., a long short-term memory (LSTM) NN [53], that maps sequences of latent states, actions, and parameters to latent next state distribution $p(\mathbf{z}_{t+1}|\mathbf{z}_{t-H:t},\mathbf{u}_{t-H:t},\mathbf{p})$. In particular, we can write the latent next states $\mathbf{z}_{t+1}$ as:
$\begin{split}z_{i,t+1}&=f_{i}^{\xi}(\mathbf{z}_{t-H:t},\mathbf{u}_{t-H:t},\mathbf{p})+\epsilon_{\xi},\quad\epsilon_{\xi}\sim\mathcal{N}(0,\sigma^{2}_{\xi})\,,\\ f_{i}^{\xi}(\bullet)&\sim\mathcal{GP}(\mu(g_{\xi}(\bullet;\boldsymbol{\theta}_{\xi}),k(g_{\xi}(\bullet;\boldsymbol{\theta}_{\xi}),g_{\xi}(\bullet^{\prime};\boldsymbol{\theta}_{\xi});\boldsymbol{\gamma}_{{\xi},i}))\,,\quad 1\leq i\leq|\mathbf{z}|\,,\end{split}$ (10)
where $z_{i,t+1}$ is sampled from the $i^{th}$ GP, $H$ is the history length, and $\epsilon_{\xi}$ is a noise term. Similarly to the encoder, the GP inputs $g_{\xi}(\mathbf{z}_{t-H:t},\mathbf{u}_{t-H:t},\mathbf{p};\boldsymbol{\theta}_{\xi})$ and $g_{\xi}(\mathbf{z}^{\prime}_{t-H:t},\mathbf{u}^{\prime}_{t-H:t},\mathbf{p}^{\prime};\boldsymbol{\theta}_{\xi})$ are the representations of obtained from the NN $g_{\xi}(\bullet;\boldsymbol{\theta}_{\xi})$ with parameters $\boldsymbol{\theta}_{\xi}$.
@@ 3.3 Training Objectives
For training the dynamical model $\xi$, we do not rely on true state values $\mathbf{s}$, but on the sequences of measurements collected at different time steps $t$. Therefore, we are not able to use supervised learning techniques to optimize the NN parameters and kernel hyperparameters. However, we can use the posterior distribution $p(\mathbf{z}_{t+1}|\mathbf{x}_{t+1})$ (see Equation (8)) as the target distribution for learning $p(\mathbf{z}_{t+1}|\mathbf{z}_{t-H:t},\mathbf{u}_{t-H:t},\mathbf{p})$. This process translates into evaluating the Kullback-Leiber (KL) divergence between these two distributions:
$\begin{split}&\mathcal{R}_{\text{reg}}^{(1)}(\boldsymbol{\theta}_{\phi},\boldsymbol{\gamma}_{\phi},\sigma^{2}_{\phi},\boldsymbol{\theta}_{\psi},\boldsymbol{\theta}_{\xi},\boldsymbol{\gamma}_{\xi},\sigma^{2}_{\xi})\\ =\leavevmode\nobreak\ &\text{KL}\left[p(\mathbf{z}_{t+1}|\mathbf{x}_{t+1})||\mathbb{E}_{\mathbf{z}_{t-H:t}|\mathbf{x}_{t-H:t}}p(\mathbf{z}_{t+1}|\mathbf{z}_{t-H:t},\mathbf{u}_{t-H:t},\mathbf{p})\right]\,,\end{split}$ (12)
where the superscript $(1)$ in $\mathcal{R}_{\text{reg}}^{(1)}$ indicates that we perform a 1-step ahead prediction of the latent states from time step $t$ to $t+1$. The distribution $p(\mathbf{z}_{t+1}|\mathbf{x}_{t+1})$ is obtained by encoding the next measurement $\mathbf{x}_{t+1}$ through $\phi$ (see Figure 2), $p(\mathbf{z}_{t+1}|\mathbf{z}_{t-H:t},\mathbf{u}_{t-H:t},\mathbf{p})$ is the distribution obtained by the forward dynamical model $\xi$ (see Figure 4), and the expectation $\mathbb{E}_{\mathbf{z}_{t-H:t}|\mathbf{x}_{t-H:t}}$ indicates the marginalization of a sequence of latent variables $\mathbf{z}_{t-H:t}$ obtained by feeding the measurements $\mathbf{x}_{t-H:t}$ to the encoder $\phi$. Different from [41], we extend this loss to $T$-step predictions:
@@
$\begin{split}&\mathcal{R}_{\text{reg}}^{(i)}(\boldsymbol{\theta}_{\phi},\boldsymbol{\gamma}_{\phi},\sigma^{2}_{\phi},\boldsymbol{\theta}_{\psi},\boldsymbol{\theta}_{\xi},\boldsymbol{\gamma}_{\xi},\sigma^{2}_{\xi})\\ =\leavevmode\nobreak\ &\text{KL}\left[p(\mathbf{z}_{t+i}|\mathbf{x}_{t+i})||\mathbb{E}_{\mathbf{z}_{t+i-1-H:t+i-1}|\mathbf{x}_{t+i-1-H:t+i-1}}p(\mathbf{z}_{t+i}|\mathbf{z}_{t+i-1-H:t+i-1},\mathbf{u}_{t+i-1-H:t+i-1},\mathbf{p})\right].\end{split}$ (14)
Similar to Equation (12), the expectation $\mathbb{E}_{\mathbf{z}_{t+i-1-H:t+i-1}|\mathbf{x}_{t+i-1-H:t+i-1}}$ indicates the marginalization of a sequence of latent variables $\mathbf{z}_{t+i-1-H:t+i-1}$ obtained by feeding the measurements $\mathbf{x}_{t+i-1-H:t+i-1}$ to the encoder $\phi$ (Figure 2).
@@
$\mathcal{R}_{\text{recon-next}}^{(i)}(\boldsymbol{\theta}_{\phi},\boldsymbol{\gamma}_{\phi},\sigma^{2}_{\phi},\boldsymbol{\theta}_{\psi},\boldsymbol{\theta}_{\xi},\boldsymbol{\gamma}_{\xi},\sigma^{2}_{\xi})=\left\|\mathbf{x}_{t+i}-\mathbb{E}[\hat{\mathbf{x}}_{t+i}|\mathbf{x}_{t+i-1-H:t+i-1}]\right\|^{2}\,,$ (16)
and the marginalized $\hat{\mathbf{x}}_{t+i}$ in the expectation $\mathbb{E}[\hat{\mathbf{x}}_{t+i}|\mathbf{x}_{t+i-1-H:t+i-1}]$ is sampled by encoding the sequence of measurements $\mathbf{x}_{t+i-1-H:t+i-1}$ through $\phi$ (see Figure 2), predicting the next latent states through $\xi$ (see Figure 4), and then decoding through $\psi$ (see Figure 3):
$\begin{split}&\mathbb{E}[\hat{\mathbf{x}}_{t+i}|\mathbf{x}_{t+i-1-H:t+i-1}]\\ =\leavevmode\nobreak\ &\mathbb{E}_{\mathbf{z}_{t+i-1-H:t+i-1}|\mathbf{x}_{t+i-1-H:t+i-1}}\mathbb{E}_{\mathbf{z}_{t+i}|\mathbf{z}_{t+i-1-H:t+i-1},\mathbf{u}_{t+i-1-H:t+i-1},\mathbf{p}}\mathbb{E}_{\hat{\mathbf{x}}_{t+i}|\mathbf{z}_{t+i}}\hat{\mathbf{x}}_{t+i}.\end{split}$ (17)
Eventually, the two SVDKL models, $\phi$ and $\xi$, utilize variational inference (VI) to approximate the posterior distributions in Equations (8) and (10), respectively. Thus, in addition to the aforementioned loss terms, we include the VI losses $\text{KL}[p(\mathbf{v})||q(\mathbf{v})]$ for both SVDKL models. Here $p(\mathbf{v})$ is the posterior to be approximated over the inducing points $\mathbf{v}$, and $q(\mathbf{v})$ represents an approximating candidate distribution. The total VI loss term is the sum of the encoder’s VI loss $\mathcal{L}_{\text{VI}}^{\phi}(\boldsymbol{\gamma}_{\phi},\sigma_{\phi}^{2})$ and the forward dynamical model’s VI loss $\mathcal{L}_{\text{VI}}^{\xi}(\boldsymbol{\gamma}_{\xi},\sigma_{\xi}^{2})$:

### v3

@@ 1 Introduction
A digital twin [1, 2, 3, 4] is a virtual replica that mimics the structure, context, and behavior of a physical system. The digital twin is synchronized to its physical twin with data in real time to predict dynamical behaviors and inform critical decisions. Constructing a digital twin typically requires accurate modeling of complex physical phenomena that are often described by nonlinear, time-dependent, parametric partial differential equations (PDEs). Full-order models (FOMs) generated by detailed solvers are often computationally demanding and thus not suitable for the multi-query, real-time contexts in digital twinning, such as uncertainty quantification [5, 6, 7], optimal control [8, 9, 10], shape optimization [11], parameter estimation [12, 13], and model calibration [14, 15]. Therefore, constructing computationally efficient reduced-order models (ROMs) with controlled accuracy is crucial for dealing with real-world systems [16].
@@
Generally speaking, any surrogate model that reduces the computational cost of a FOM can be considered a ROM. ROMs aim to intelligently represent high-dimensional dynamical systems in carefully-established low-dimensional latent spaces, so that the low-dimensionality improves computational efficiency, and the accuracy remains under control [17, 18, 19, 20, 21, 22, 16]. Despite the wide variety of approaches that can be found in the literature, we can identify two major categories of ROM approaches: intrusive and non-intrusive methods. Intrusive ROM techniques require access to the full-order solvers, which is often inconvenient in industrial implementations, especially for the cases involving legacy code and/or readily executed software. To overcome this limitation, non-intrusive ROM techniques have been developed to learn low-dimensional representations primarily from data, without accessing the code of FOMs. Such flexibility of non-intrusive ROMs has recently motivated the development of a vast amount of data-driven methods, such as dynamic mode decomposition [23, 24], reduced-order operator inference [25, 26, 27, 28], sparse identification of reduced latent dynamics [29, 30, 31, 32, 33, 34], manifold learning using deep auto-encoders [35, 36, 37, 38], data-driven approximation of time-integration schemes [39], Gaussian process modeling for low-dimensional representations [40, 41], and kernel flows [42].
@@
However, learning ROMs from noisy data is extremely challenging. While many works in the literature have focused on learning compact representation of high-dimensional data, for example using neural networks (NNs) [43, 44], or quantifying uncertainties using, for example, Gaussian processes (GPs) [45], herein we argue that both challenges must be tackled simultaneously for discovering ROMs properly. NNs excel in learning complex and nonlinear dependencies of the data, while they tend to struggle when quantifying uncertainties. Conversely, GPs struggle with large datasets due to the limited expressivity of the kernels and the need for the inversion of the covariance matrix, while they excel to quantify uncertainties. To overcome these limitations, and to leverage on the positive features of both NNs and GPs synergistically, deep kernel learning (DKL) and stochastic variational DKL (SVDKL) were introduced in recent years [46, 47]. DKL aims to combine the best of the GP and NN worlds by constructing expressive deep kernels that can model complex relations of the data. DKL builds a deep kernel by feeding to a GP the data processed by a NN. In addition, SVDKL relies of variational inference [48] to allow for the use of traditional minibatch training techniques employed by NN-based models, making SVDKL suitable for effectively dealing with large dataset. Variational inference amortizes the cost of sampling from the (non-Gaussian) posterior distribution by approximating the distribution with the best-fitting Gaussian, reducing the computational cost of SVDKL models.
Another approach to learn expressive deep kernels is kernel flows [42, 49, 50]. Kernel flows are based on the simple idea that good kernels maintain similar accuracy if the data points are reduced. Kernel flows progressively refine the kernel using subsets of the training set. Thus, they offer a solution to scale GPs to large datasets. Similarly to DKL, kernel flows can use sequences of nonlinear transformations to pre-process the data before feeding them to the GP kernel. Kernel flows have been successfully used for modeling low-dimensional dynamical systems [51, 52, 53] even in the case of irregular sampling of the data. Although kernel flows have been applied to low-dimensional dynamical systems, it is worth mentioning that this approach is not inherently limited by the data dimensionality. For example, kernel flows have also been successfully applied to the MNIST dataset [42].
@@
In this paper, we exploit the SVDKL idea to develop a non-intrusive ROM that can deal with both high-dimensional and noisy data. In particular, with reference to Figure 1, our method includes a SVDKL encoder to compress high-dimensional measurements into low-dimensional distributions of state variables, a recurrent SVDKL latent dynamical model to predict the system’s evolution over time, and a decoder to reconstruct the measurements and interpret the latent representations.
@@
Figure 1: Proposed framework for reduced-order modeling of dynamical systems. We encode the measurements $\mathbf{x}_{t}$ at different time instances into the latent variables $\mathbf{z}_{t}$ by means of a deep kernel learning encoder. Then, we feed a sequence of length $H$ of consecutive latent variables $\mathbf{z}_{t-H:t}$, actions $\mathbf{u}_{t-H:t}$, and parameters $\mathbf{p}$ to a recurrent deep kernel learning to predict the next latent variable $\mathbf{z}_{t+1}$. The measurements are then reconstructed $\hat{\mathbf{x}}$ by means of a decoder from the latent variables.
@@
The model is trained without labeled data in an unsupervised manner by only relying on high-dimensional and noisy measurements of the system. We show the capabilities of our framework in two challenging numerical examples, namely, the dynamics of (i) a double pendulum and of (ii) a distributed reaction-diffusion system, assuming in both cases to deal with measurements over a two-dimensional spatial region at different time instants – thus mimicking a set of observations acquired by a camera. Our contribution is threefold:
@@
  • we improve the prediction accuracy and consistency over long horizons of the SVDKL framework introduced in [41] by modeling the latent dynamics using a recurrent NN,
@@
The paper is organized as follows: in Section 2, we introduce the building blocks of our framework, namely GPs, and DKL, and in Section 3, we describe our novel DKL-based method. Section 4 presents the numerical experiments, shows and discusses the results, and Section 5 concludes the paper.
@@ 2.2 Deep Kernel Learning
The GP hyperparameters $\boldsymbol{\gamma}$ and the NN parameters $\boldsymbol{\theta}$ are jointly learned by maximizing the log marginal likelihood (see Equation (3)). It is possible to use the chain rule to compute derivatives of the log marginal likelihood, that we indicate with $\mathcal{L}(\boldsymbol{\gamma},\boldsymbol{\theta})$, with respect to kernel hyperparameters $\boldsymbol{\gamma}$ and the NN parameters $\boldsymbol{\theta}$, thus obtaining:
@@
Optimizing the GP hyperparameters through Equation (3) requires to repeatedly invert the covariance matrix $k_{\boldsymbol{\gamma}}(X,X)+\sigma_{\epsilon}^{2}I$, with size $N\times N$. However, for large datasets, i.e., when $N\gg 1$, the full covariance matrix may be too large to be stored in memory and computationally prohibitive to invert it. In these cases, it is common practice in machine learning to utilize minibatches of randomly-samples data points to compute the loss functions and compute the gradients. However, when we use minibatches to train the GPs, the posterior distribution is not Gaussian anymore, even when a Gaussian likelihood is used. When dealing with non-Gaussian posteriors, we can rely on variational inference (VI) [48]. VI is an approximation technique for dealing with non-Gaussian posteriors, deriving from minibatch training and/or non-Gaussian likelihoods. VI allows one to approximate the posterior by the best-fitting Gaussian distribution using a set of samples, often called inducing points, from the posterior. VI allows for drastically reducing the computation cost of standard GPs and effectively utilizing DKL with large datasets.
@@ 3.1 Problem Statement
For conciseness, we use simplified notation $\mathbf{x}_{t}^{-H}$ to represent the series $\mathbf{x}_{t-H},\cdots,\mathbf{x}_{t}$ throughout the paper, with $H$ denoting the history length. This applies not only to the full states $\mathbf{x}$ but also to the latent states $\mathbf{z}$ and control inputs $\mathbf{u}$.
@@ 3.2 Model Architecture
    a recurrent SVDKL forward dynamical model predicting the evolution of the system using a history of latent state variables $\mathbf{z}_{t}^{-H}$ and control inputs $\mathbf{u}_{t}^{-H}$, and parameters $\mathbf{p}$, if available.
@@
With reference to Figure 3, the decoder $\psi$ is modeled using an NN with parameters $\boldsymbol{\theta}_{\psi}$. The decoder maps a sample $\mathbf{z}$ to the conditional distribution $p(\hat{\mathbf{x}}|\mathbf{z})$. Similar to VAEs [56], the distribution $p(\hat{\mathbf{x}}|\mathbf{z})$ is approximated to be independently-distributed Gaussian distributions. In general, the decoding NN $\psi$ learns the means and variances of these distributions. Following [57, 58], however, our decoder in this work only learns the mean vector of $\hat{\mathbf{x}}|\mathbf{z}$, while the covariance is set to the identity matrix $I$, i.e., $\psi:\mathbb{R}^{|\mathbf{z}|}\rightarrow[0,1]^{|\mathbf{x}|}$ and
@@
With reference to Figure 4, the forward dynamical model $\xi:\mathbb{R}^{|\mathbf{z}|}\times\mathbb{R}^{|\mathbf{u}|}\times\mathbb{R}^{|\mathbf{p}|}\rightarrow[0,1]^{|\mathbf{z}|}$, with parameters $\boldsymbol{\theta}_{\xi}$ and hyperparameters $\boldsymbol{\gamma}_{\xi},\sigma_{\xi}^{2}$, is a SVDKL with a recurrent NN architecture, i.e., a long short-term memory (LSTM) NN [59], that maps sequences of latent states, actions, and parameters to latent next state distribution $p\left(\mathbf{z}_{t+1}|\mathbf{z}_{t}^{-H},\mathbf{u}_{t}^{-H},\mathbf{p}\right)$.
In particular, we can write the latent next states $\mathbf{z}_{t+1}$ as:
$\begin{split}z_{i,t+1}&=f_{i}^{\xi}\left(\mathbf{z}_{t}^{-H},\mathbf{u}_{t}^{-H},\mathbf{p}\right)+\epsilon_{\xi},\quad\epsilon_{\xi}\sim\mathcal{N}(0,\sigma^{2}_{\xi})\,,\\ f_{i}^{\xi}(\bullet)&\sim\mathcal{GP}(\mu(g_{\xi}(\bullet;\boldsymbol{\theta}_{\xi}),k(g_{\xi}(\bullet;\boldsymbol{\theta}_{\xi}),g_{\xi}(\bullet^{\prime};\boldsymbol{\theta}_{\xi});\boldsymbol{\gamma}_{{\xi},i}))\,,\quad 1\leq i\leq|\mathbf{z}|\,,\end{split}$ (10)
where $z_{i,t+1}$ is sampled from the $i^{th}$ GP, $H$ is the history length, and $\epsilon_{\xi}$ is a noise term. Similarly to the encoder, the GP inputs $g_{\xi}\left(\mathbf{z}_{t}^{-H},\mathbf{u}_{t}^{-H},\mathbf{p};\boldsymbol{\theta}_{\xi}\right)$ and $g_{\xi}\left((\mathbf{z}^{\prime})_{t}^{-H},(\mathbf{u}^{\prime})_{t}^{-H},\mathbf{p}^{\prime};\boldsymbol{\theta}_{\xi}\right)$ are the representations of obtained from the NN $g_{\xi}(\bullet;\boldsymbol{\theta}_{\xi})$ with parameters $\boldsymbol{\theta}_{\xi}$.
@@ 3.3 Training Objectives
For training the dynamical model $\xi$, we do not rely on true state values $\mathbf{s}$, but on the sequences of measurements collected at different time steps $t$. Therefore, we are not able to use supervised learning techniques to optimize the NN parameters and kernel hyperparameters. However, we can use the posterior distribution $p(\mathbf{z}_{t+1}|\mathbf{x}_{t+1})$ (see Equation (8)) as the target distribution for learning $p\left(\mathbf{z}_{t+1}|\mathbf{z}_{t}^{-H},\mathbf{u}_{t}^{-H},\mathbf{p}\right)$. This process translates into evaluating the Kullback-Leiber (KL) divergence between these two distributions:
$\begin{split}&\mathcal{R}_{\text{reg}}^{(1)}(\boldsymbol{\theta}_{\phi},\boldsymbol{\gamma}_{\phi},\sigma^{2}_{\phi},\boldsymbol{\theta}_{\psi},\boldsymbol{\theta}_{\xi},\boldsymbol{\gamma}_{\xi},\sigma^{2}_{\xi})\\ =\leavevmode\nobreak\ &\text{KL}\left[p(\mathbf{z}_{t+1}|\mathbf{x}_{t+1})\Big{\|}\leavevmode\nobreak\ \mathbb{E}_{\mathbf{z}_{t}^{-H}|\mathbf{x}_{t}^{-H}}\leavevmode\nobreak\ p\left(\mathbf{z}_{t+1}|\mathbf{z}_{t}^{-H},\mathbf{u}_{t}^{-H},\mathbf{p}\right)\right]\,,\end{split}$ (12)
where the superscript $(1)$ in $\mathcal{R}_{\text{reg}}^{(1)}$ indicates that we perform a 1-step ahead prediction of the latent states from time step $t$ to $t+1$ and the subscript reg stands for regularization. The distribution $p(\mathbf{z}_{t+1}|\mathbf{x}_{t+1})$ is obtained by encoding the next measurement $\mathbf{x}_{t+1}$ through $\phi$ (see Figure 2), $p\left(\mathbf{z}_{t+1}|\mathbf{z}_{t}^{-H},\mathbf{u}_{t}^{-H},\mathbf{p}\right)$ is the distribution obtained by the forward dynamical model $\xi$ (see Figure 4), and the expectation $\mathbb{E}_{\mathbf{z}_{t}^{-H}|\mathbf{x}_{t}^{-H}}$ indicates the marginalization of a sequence of latent variables $\mathbf{z}_{t}^{-H}$ obtained by feeding the measurements $\mathbf{x}_{t}^{-H}$ to the encoder $\phi$. Different from [41], we extend this loss to $T$-step predictions:
@@
$\begin{split}&\mathcal{R}_{\text{reg}}^{(i)}(\boldsymbol{\theta}_{\phi},\boldsymbol{\gamma}_{\phi},\sigma^{2}_{\phi},\boldsymbol{\theta}_{\psi},\boldsymbol{\theta}_{\xi},\boldsymbol{\gamma}_{\xi},\sigma^{2}_{\xi})\\ =\leavevmode\nobreak\ &\text{KL}\left[p(\mathbf{z}_{t+i}|\mathbf{x}_{t+i})\Big{\|}\leavevmode\nobreak\ \mathbb{E}_{\mathbf{z}_{t+i-1}^{-H}\big{|}\mathbf{x}_{t+i-1}^{-H}}p\left(\mathbf{z}_{t+i}\leavevmode\nobreak\ |\leavevmode\nobreak\ \mathbf{z}_{t+i-1}^{-H},\mathbf{u}_{t+i-1}^{-H},\mathbf{p}\right)\right].\end{split}$ (14)
Similar to Equation (12), the expectation $\mathbb{E}_{\mathbf{z}_{t+i-1}^{-H}|\mathbf{x}_{t+i-1}^{-H}}$ indicates the marginalization of a sequence of latent variables $\mathbf{z}_{t+i-1}^{-H}$ obtained by feeding the measurements $\mathbf{x}_{t+i-1}^{-H}$ to the encoder $\phi$ (Figure 2).
@@
$\mathcal{R}_{\text{recon-next}}^{(i)}(\boldsymbol{\theta}_{\phi},\boldsymbol{\gamma}_{\phi},\sigma^{2}_{\phi},\boldsymbol{\theta}_{\psi},\boldsymbol{\theta}_{\xi},\boldsymbol{\gamma}_{\xi},\sigma^{2}_{\xi})=\left\|\mathbf{x}_{t+i}-\mathbb{E}[\hat{\mathbf{x}}_{t+i}\leavevmode\nobreak\ |\leavevmode\nobreak\ \mathbf{x}_{t+i-1}^{-H},\mathbf{u}_{t+i-1}^{-H},\mathbf{p}]\right\|^{2}\,,$ (16)
and the marginalized $\hat{\mathbf{x}}_{t+i}$ in the expectation $\mathbb{E}\left[\hat{\mathbf{x}}_{t+i}|\mathbf{x}_{t+i-1}^{-H}\right]$ is sampled by encoding the sequence of measurements $\mathbf{x}_{t+i-1}^{-H}$ through $\phi$ (see Figure 2), predicting the next latent states through $\xi$ (see Figure 4), and then decoding through $\psi$ (see Figure 3):
$\mathbb{E}\left[\hat{\mathbf{x}}_{t+i}\leavevmode\nobreak\ |\leavevmode\nobreak\ \mathbf{x}_{t+i-1}^{-H},\mathbf{u}_{t+i-1}^{-H},\mathbf{p}\right]=\mathbb{E}_{\mathbf{z}_{t+i-1}^{-H}|\mathbf{x}_{t+i-1}^{-H}}\mathbb{E}_{\mathbf{z}_{t+i}|\mathbf{z}_{t+i-1}^{-H},\mathbf{u}_{t+i-1}^{-H},\mathbf{p}}\mathbb{E}\left[\hat{\mathbf{x}}_{t+i}|\mathbf{z}_{t+i}\right].$ (17)
Eventually, the two SVDKL models, $\phi$ and $\xi$, utilize VI to approximate the posterior distributions in Equations (8) and (10), respectively. Thus, in addition to the aforementioned loss terms, we include the VI losses $\text{KL}[p(\mathbf{v})||q(\mathbf{v})]$ for both SVDKL models. Here $p(\mathbf{v})$ is the posterior to be approximated over the inducing points $\mathbf{v}$, and $q(\mathbf{v})$ represents an approximating candidate distribution. The total VI loss term is the sum of the encoder’s VI loss $\mathcal{L}_{\text{VI}}^{\phi}(\boldsymbol{\gamma}_{\phi},\sigma_{\phi}^{2})$ and the forward dynamical model’s VI loss $\mathcal{L}_{\text{VI}}^{\xi}(\boldsymbol{\gamma}_{\xi},\sigma_{\xi}^{2})$: